# Normal forms for ordinary differential operators, I

### v1

@@ 4.3 Commutativity criterion for normal forms having the restriction top line
By the same arguments as above (use Lemma 5.3 item 1) there doesn’t exist $(l,j)\in E(L_{1}M_{3})\bigcup E(L_{3}M_{1})\bigcup E(L_{3}M_{3})$, such that
@@
$H_{N_{F}}(F(P^{\prime},Q^{\prime}))=\\ \sum_{l=0}^{s-1}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})+\sum_{l=s}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})\\ =\sum_{l=s}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})$
@@
$HS_{N_{F}}^{sa_{0}}(F(P^{\prime},Q^{\prime}))=HS_{N_{F}}^{sa_{0}}(H_{N_{F}}(F(P^{\prime},Q^{\prime})))\\ =\sum_{j=1}^{m}\binom{u_{j}}{s}k_{j}\cdot HS_{N_{F}}^{sa_{0}}(\sum_{h=0}^{s}\binom{s}{h}\mathbb{L}_{0}^{s-h}\mathbb{L}_{1}^{h}\partial^{N_{F}-sp})+\\ \sum_{l=s+1}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot HS_{N_{F}}^{sa_{0}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})\\ =\sum_{j=1}^{m}\binom{u_{j}}{s}k_{j}\mathbb{L}_{0}^{s}\partial^{N_{F}-sp}$
@@ 1. Obviously, for any operator $P$ from formulation we have
We now want to estimate $v_{\sigma,\rho}(LM)$ and $v_{\sigma,\rho}([L,M])$ with the help of $v_{\sigma,\rho}(L)$ and $v_{\sigma,\rho}(M)$ (cf. similar estimations for $L,M\in A_{1}$ in [10, L.2.7]). We consider first the case when $L,M$ are monomials from $Hcpc(k)$.
@@ 1. Suppose
  3. (c)
    $f_{\sigma,\rho}(L)=0,f_{\sigma,\rho}(M)\neq 0$. This case is analogous to b), so we omit the details.
  4. (d)
    $f_{\sigma,\rho}(L)=0,f_{\sigma,\rho}(M)=0$, in this case just deal with $L,M$ like in (b), the discussion will be the same, we omit the details.

### v3

@@ 4.3 Commutativity criterion for normal forms having the restriction top line
By the same arguments as above (use Lemma 5.3 item 1) there doesn’t exist
$(l,j)\in E(L_{1}M_{3})\bigcup E(L_{3}M_{1})\bigcup E(L_{3}M_{3})$, such that
@@
$H_{N_{F}}(F(P^{\prime},Q^{\prime}))=\sum_{l=0}^{s-1}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})+\sum_{l=s}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})\\ =\sum_{l=s}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})$
@@
$HS_{N_{F}}^{sa_{0}}(F(P^{\prime},Q^{\prime}))=HS_{N_{F}}^{sa_{0}}(H_{N_{F}}(F(P^{\prime},Q^{\prime})))=\sum_{j=1}^{m}\binom{u_{j}}{s}k_{j}\cdot HS_{N_{F}}^{sa_{0}}(\sum_{h=0}^{s}\binom{s}{h}\mathbb{L}_{0}^{s-h}\mathbb{L}_{1}^{h}\partial^{N_{F}-sp})+\\ \sum_{l=s+1}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot HS_{N_{F}}^{sa_{0}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})=\sum_{j=1}^{m}\binom{u_{j}}{s}k_{j}\mathbb{L}_{0}^{s}\partial^{N_{F}-sp}$
@@ 1. Obviously, for any operator $P$ from formulation we have
We now want to estimate $v_{\sigma,\rho}(LM)$ and $v_{\sigma,\rho}([L,M])$ with the help of $v_{\sigma,\rho}(L)$ and $v_{\sigma,\rho}(M)$ (cf. similar estimations for $L,M\in A_{1}$ in [14, L.2.7]). We consider first the case when $L,M$ are monomials from $Hcpc(k)$.
@@ 1. Suppose
(c) $f_{\sigma,\rho}(L)=0,f_{\sigma,\rho}(M)\neq 0$. This case is analogous to b), so we omit the details.
(d) $f_{\sigma,\rho}(L)=0,f_{\sigma,\rho}(M)=0$, in this case just deal with $L,M$ like in (b), the discussion will be the same, we omit the details.